OpenStudy (anonymous):

if the measure of an angle is 90-x what is the measure of its complement?

OpenStudy (unklerhaukus):

x

OpenStudy (anonymous):

seriously? dont waste my time and not answer the question!

jimthompson5910 (jim_thompson5910):

Original angle is 90-x Measure of complement angle is 90-(90-x) = 90-90+x = 0+x = x So the complement is x degrees So the answer is simply x

OpenStudy (unklerhaukus):

ok, two angles are said to complementry if they sum to a right angle i.e. if a+b = 90 (a and b are complementary) so Your question is asking 90-x + a= 90

OpenStudy (unklerhaukus):

a=x